OpenStudy (anonymous):

1.Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

$0.50 to purchase sale price = .50(1.40) = $ 0.70 this means they make $ 0.20 profit per flower 0.20f = 500 f = 500/0.20 f = 2500 Angelica's service club would have to sell 2500 flowers to clear a profit of $500. ==================================== 10A + 6C = 80 C = 5 now sub 10A + 6(5) = 80 10A + 30 = 80 10A = 80 - 30 10A = 50 A = 5 check.. 10A + 6C = 80 10(5) + 6(5) = 80 50 + 30 = 80 80 = 80 (correct) If there were 5 children in the group, that means there were also 5 adults. =======================================... | x - 0.525| = 0.01 x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 -| x - 0.525| = 0.01 -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515 minimum is : -0.515 diameter maximum is : 0.535 diameter